OpenStudy (anonymous):

The perimeter of a rectangle is 54 inches, and its area is 180 in^2. Find its length and width.

OpenStudy (campbell_st):

so for perimeter P = 2l + 2w or 54 = 2l + 2w (equation 1) next for the area A= lw or 180 = lw (equation 2) rewrite equation 2 and 27 - w = l and substitute into equation 2 180 =(27 - w)w so you now have a quadratic that needs to be solved \[-(w^2 - 27w + 180) = 0\] solve for the w, but only use the positive answers, then substitute into equation 1 to find l.